A hemispherical container contains water to a depth of 10 cm. If the volume of the water inside the container is 5000m/3 cubic centimeter, find the diameter of the vessel.


1
Expert's answer
2021-11-25T23:37:44-0500


The volume of water = (5000 *"\pi")/3 cm3

The water is contained to a depth of 10 cm

Let radius of hemisphere be R1 cm

Let radius of the water be R2 cm


Now volume of water = (2/3)*("\pi")*(radius)3

(5000 * "\pi")/3 = (2/3)*("\pi")*(R2)3

(R2)3 = 5000/2 = 2500 => R2 = cube root (2500) = 13.5 cm


Now as per the diagram, AC = AD

AB + BC = R1 (Since AD = radius of hemisphere which is R1)

10 + R2 = R1

R1 =10 + 13.5 = 23.5 cm

So,

diameter of hemisphere = 2 * radius = 2 * 23.5 = 47 cm